In the RLC Parallel circuit , Is=?, L=7 H, R=6 and C=(1/42) F. IR=? IL(T)=? IC=?
in the drawing (I(s) next R next L final C )
1
Expert's answer
2020-05-08T14:12:57-0400


Assume that the voltage source produces the voltage of RMS value of 220 V and frequency 50 Hz. The reactances of the inductance and capacitance are


"X_L=2\pi fL,\\
X_C=\frac{1}{2\pi fC}."

Since this is a parallel connection, apply Ohm's law to each element to find current:


"I_R=\frac{V}{R}=\frac{220}{6}=36.7\text{ A}.
\\
I_L=\frac{V}{X_L}=\frac{220}{2\pi\cdot50\cdot7}=0.1\text{ A}.
\\
I_C=\frac{V}{X_C}=\frac{220\cdot42}{2\pi\cdot 50}=29.4\text{ A}."

The source current is


"I_s=\sqrt{I_R^2+(I_L-I_C)^2}=22\text{ A}."